Sehro:

A mixture placed in an Erlenmeyer flask comprises 6mL of silica gel and 40mL of a solvent containing, in solution, 100 mg of a non-volatile compound. After stirring, the mixture was left to stand before a 10mL aliquot of the solution was extracted and evaporated to dryness. The residue weighed 12 mg

Gdeinward:

Hello, hope you are doing well. In the future, please include the question in its entirety. You cant answer an incomplete statement. Below is the full question: A mixture placed in an Erlenmeyer flask comprises 6 mL of silica gel and 40 mL of a solvent containing, in solution, 100 mg of a non-volatile compound. After stirring, the mixture was left to stand before a 10 mL aliquot of the solution was extracted and evaporated to dryness. The residue weighed 12 mg. Calculate the adsorption coefficient, K = CS/CM, of the compound in this experiment.

Gdeinward:

@vocaloid

Vocaloid:

the silica gel is the stationary phase and the solvent is the mobile phase. K = CS/CSM = (molar concentration of solute in stationary phase) / (molar conc. in mobile phase) the same 100 mg compound will be distributed in some proportion between the mobile and stationary phases, meaning we can simply calculate how many grams are in the stationary vs. mobile phase. let's look at the 10 mL aliquot. 12 mg of the compound in 10 mL solution --> 12mg/10mL in the mobile phase. so in the total 40 mL of the mobile phase, there are 48mg, and the proportion is 48mg/40mL for the mobile phase. This is your CM. Repeat this process for the stationary phase. The amount of compound left in the stationary phase is 100mg minus the amount in the mobile phase. Divide by the volume of the stationary phase. This is your CS. Finally calculate K = CS/CM